OpenStudy (anonymous):

What is the solution to the system? Line 1 x y –1 9 3 –3 Line 2 x y –1 –1 2 5

OpenStudy (anonymous):

Can someboy help please

OpenStudy (anonymous):

opps sombody

OpenStudy (anonymous):

@pounce0129

OpenStudy (pounce0129):

i did this 2 years ago, so im trying to think of a way to solve it

OpenStudy (anonymous):

ok thank you

OpenStudy (pounce0129):

do you just need the equations of both lines or to solve the systems?

OpenStudy (pounce0129):

@thedomination100

OpenStudy (anonymous):

it says this

OpenStudy (pounce0129):

ok i got the equations of the lines, do you think you can take it from there?

OpenStudy (anonymous):

ok

OpenStudy (pounce0129):

so the equation of line 1 is: y=-3x+6 equation of line 2 is y=3x-1

OpenStudy (pounce0129):

so then, from there you just use the elimination method

OpenStudy (anonymous):

i still dont get it

OpenStudy (pounce0129):

x=2, y=5

OpenStudy (pounce0129):

do you want me to explain how got it?

OpenStudy (anonymous):

yes please

OpenStudy (pounce0129):

ok, so you find the slope of the first line, then you insert that value into the slope intercept form

OpenStudy (pounce0129):

so for the first one \[\frac{ -3-9 }{ 3-(-1) } = \frac{ -12 }{ 4}=-3\]

OpenStudy (pounce0129):

so -3 is your slope for the first line

OpenStudy (pounce0129):

then you use \[y=mx+b\] and insert a point and the slope

OpenStudy (pounce0129):

do you get it so far?

OpenStudy (anonymous):

yes thank you

OpenStudy (pounce0129):

so then your equation would be \[9=-3(-1)+b\] and you solve for b

OpenStudy (pounce0129):

so then the slope intercept equation for the first line would be y=-3x+6

OpenStudy (pounce0129):

and you do the same thing or the other line

OpenStudy (pounce0129):

understand?

OpenStudy (pounce0129):

well i gotta go